8e^(x+4)=8 write in terms of logs x=?
Answer by lwsshak3(11628)

8e^(x+4)=8 write in terms of logs x=?
8e^(x+4)=8
divide both sides by 8
e^(x+4)=1
take log(ln) of both sides
(x+4) lne=ln(1)
log of any base=1
log of (1) to any base=0
x+4=0
x=-4